# A Hyperparameter Study for Quantum Kernel Methods

### v2

@@ Abstract
Quantum kernel methods are a promising method in quantum machine learning thanks to the guarantees connected to them. Their accessibility for analytic considerations also opens up the possibility of prescreening datasets based on their potential for a quantum advantage. To do so, earlier works developed the geometric difference, which can be understood as a closeness measure between two kernel-based machine learning approaches, most importantly between a quantum kernel and classical kernel. This metric links the quantum and classical model complexities. Therefore, it raises the question of whether the geometric difference, based on its relation to model complexity, can be a useful tool in evaluations other than for the potential for quantum advantage. In this work, we investigate the effects of hyperparameter choice on the model performance and the generalization gap between classical and quantum kernels. The importance of hyperparameter optimization is well known also for classical machine learning. Especially for the quantum Hamiltonian evolution feature map, the scaling of the input data has been shown to be crucial. However, there are additional parameters left to be optimized, like the best number of qubits to trace out before computing a projected quantum kernel. We investigate the influence of these hyperparameters and compare the classically reliable method of cross validation with the method of choosing based on the geometric difference. Based on the thorough investigation of the hyperparameters across 11 datasets we identified commodities that can be exploited when examining a new dataset. In addition, our findings contribute to better understanding of the applicability of the geometric difference.
@@ I Introduction
Assessing the performance of quantum models on current quantum computers in tackling machine learning tasks on classical datasets is an enduring ambition of the research community. Many strides have been made addressing this subject from theory-driven approaches that showed encouraging results for specific models, e.g. investigating the difference in capacity between classical and quantum models [2] and showcasing a model that allows to sample from a circuit that is hard to simulate classically [5], to more empirical-driven projects that showed promising results for specific models on specific use-cases, e.g. medical classification with hybrid quantum-classical convolutional neural network [15] and generation of high-resolution MNIST datapoints with hybrid generative neural networks [21]. However, making an overall claim of quantum advantage for a broader class of models for a broader set of use-cases remains challenging, especially due to the notorious insufficiency of theoretical understanding of certain sub-fields of machine learning, such as deep learning (DL) [12]. Nevertheless, in the recent years a body of literature emerged linking DL models to kernel methods (KMs) [6, 9, 11] followed by literature making similar conclusion for the quantum domain [20, 8, 22]. This is great news due to favorable qualities of KMs, such as exhibiting convex loss landscapes, which allows to establish convergence, and hence trainability, guarantees [22]. It is a big advantage in comparison to DL that infamously suffer from non-convex loss landscapes and their quantum counterparts that suffer from barren plateaus [14]. Focusing on KMs therefore allows for a more theoretically sound argument, while at the same time the insights gained for KMs can be applied to a broader range of models due to the aforementioned interrelationships.
@@
In this work, we study how the choice of hyperparameters influence the empirical performance of the quantum KMs and how it influences the GD [7]. The motivation behind this work is to pinpoint the parameters and ranges which are most crucial to optimize. A second incentive is to deepen our understanding of the relevance of the GD-tool in an empirical setup. We performed an experiment spanning multiple different datasets (classical and artificially labeled), a large hyperparameter search grid and different kernel setups. Our results indicate that the GD depends on the individual hyperparameters to a similar extent as the accuracy, but the optimal values of them differ greatly between the metrics. We found also empirically which parameters are most crucial to optimize and highlight how.
@@ II-B Hamiltonian evolution feature map
For a quantum algorithm to operate on a classical data, the data has to be embedded into a quantum state $\phi:\mathbf{x}_{i}\to\ket{\mathbf{x}_{i}}$, which exists in a complex Hilbert space $\mathcal{H}$ that is $\mathbb{C}^{2^{n}}$ with $n$ being the number of qubits. The significance of the choice of embedding operation for the performance of quantum models has been highlighted in the literature [2, 23]. In this work, we concentrate on the Hamiltonian evolution feature map as it has been proven to be powerful in QKM setting [4, 7, 24]. This feature map has been developed for many-body problems and is defined as follows
@@
where $H_{j}^{XYZ}=(X_{j}X_{j+1}+Y_{j}Y_{j+1}+Z_{j}Z_{j+1})$ with $X_{j},Y_{j},Z_{j}$ being Pauli operators acting on $j$-th qubit, |$\ket{\psi_{j}}$ is a Haar-random state, $t$ and $T$ are hyperparameters that symbolizes total evolution time and number of Trotter steps respectively. The significance of hyperparameter $t$ for QKMs has been highlighted by Shaydulin et al. [24]. Our results support and extend their work.
@@ II-C Quantum kernels
Quantum kernels insert data into exponentially high dimensional Hilbert feature space, which is intractable for classical computers. The high dimensionality of this space has to be treated with caution as it requires $N\geq\Omega(2^{n})$ of training data to approximate a certain function well, where $n$ is amount of logical qubits [7, 10, 19, 26]. This is due to a problem known as exponential concentration [26]. The same logic holds for a quantum fidelity kernel of mixed quantum states $k(x,x^{\prime})=Tr(\rho(x)\rho(x^{\prime}))$ [19]. Reducing the dimensionality of the space by projecting it onto a smaller subspace, thereby creating so-called projected quantum kernels, increases the generalization capacity of these methods. To achieve that we can utilize reduced density matrices (RDMs) $\rho^{(q)}(x)=Tr_{\neq q}(\Phi(\rho(x)))$, where $Tr_{\neq q}$ stands for a trace of a subsystem that is created by excluding $q$ qubits and $\Phi$ is a trace-preserving map. Depending on the choice of the map $\Phi$, different proposals for projected kernels can be found in the literature [7, 10]. As generalized by Nakaji et al. [19], these projected (RDM-based) kernel can be categorized into inner product based defined as
@@
In this paper an adaptation of the inner product based kernel is considered. The motivation is that the kernel entries can become relatively small for large dimensions and highly entangled states. On the main diagonal the values are related to the purity of a quantum state $Tr(\rho^{2})$, which can become $\frac{1}{d}$ where $d$ is the dimension of the Hilbert space. The normalized inner product based kernel is defined as
and therefore remains a valid kernel, because the same feature vectors as in Nakaji et al. [19] can be found for the redefined density matrices (DMs). This normalization affects all entries in the gram matrix, in particular the main diagonal now consists of just ones, as it is common for kernels in general.
@@ II-D Geometric difference (GD)
Huang et al. [7] challenged a common argument for claiming quantum advantage that is often used in QML literature. In their work, they showed that a classical intractability of the function $f(x)$, that is represented by a quantum model, is not sufficient to claim a quantum advantage. Under certain circumstances, a classical learner $h(x)$ can learn to approximate $f(x)$ given a sufficient number $N$ of independent samples from a data distribution $\mathcal{D}$. They showed that an expected error between $f(x)$ and $h(x)$ can be upper bounded as:
@@
where $s_{K}(N)$ is the model complexity of the trained function $h(x)$ and $c>0$ is a constant. For that reason, with $N\propto s_{K}(N)/\epsilon^{2}$ datapoints a classical model can learn to predict the quantum function $f(x)$ up to an error $\epsilon$. Thus, if the best classical learner $h(x)$ has a low model complexity $s_{K}(N)$, it will not require too much data $N$ to learn the outputs of the quantum model $f(x)$. From here it follows, that for a quantum advantage to exist it is required that there is the largest possible separation between complexities of quantum and classical models. Huang et al. [7] considered kernel models in their work and defined an asymmetric GD-based test between kernel matrices that captures this separation on a specific dataset. This test is defined as follows:
@@
where $K_{C}$ and $K_{Q}$ are classical and quantum kernel matrices with $Tr(K_{Q})=Tr(K_{C})=N$, respectively. The distance based and the normalized inner product based kernels meet this requirement, while the inner product based kernel requires rescaling according to $\tilde{K}_{Q}=N\cdot K_{Q}/Tr(K_{Q})$ in the case of projected kernels, due to the problem discussed in the previous section that motivated the normalized version. Importantly, this test is reliant on the GD between embedded datapoints, without considering the labeling function. The GD now links the classical model complexity $s_{K_{C}}(N)$ and the quantum model complexity $s_{K_{Q}}(N)$ according to
@@ III Experimental setup
In this work we investigate (1) how different parameter settings affect the empirical performance of quantum kernels, as well as their generalization ability, as determined by the GD; (2) whether the GD is a reliable indicator that can be used to optimize the hyperparameters; (3) if normalization can elevate the inner product based kernel. To shed some light on these points we have designed an experiment to evaluate a variety of different hyperparameter settings. In this section, we describe the parameters of the search grid used in our experiment. We varied the types of models and their hyperparameters and tested their performance on various datasets.
@@ III-A Models
In this work, we combine classical SVMs (see Section II) from sklearn library with quantum kernels implemented with PennyLane framework on simulators. To obtain the kernel, we first embed classical datapoints with the Hamiltonian evolution feature map as in Eq. 2 and the full DM of the resulting quantum state is saved. These DMs are used to compute RDM kernels described in Eq. 4 (with additional rescaling to $Tr(K_{Q})=N$), Eq. 5 and Eq. 6. From here forth we will refer to these kernels as inner, distance and inner normalized kernels respectively. In all cases, we set $\alpha_{k}=1/N_{K}$. In Appendix E we discuss the case $\alpha_{k}=1$.
@@ III-C Datasets
To train and test our models, we selected the 11 different datasets listed in Table II that have been previously used in the QML literature [7, 16, 17, 24]. Due to simulation time restrictions, we compress each dataset’s feature space to 5 dimensions, which translates to 6 qubits according to Eq. 2, and 200 datapoints (or less features/datapoints if fewer were available). In this study, we restrict to binary classification tasks, so for multi-class dataset 8 (Fashion-MNIST) only two classes are selected (T-shirt/Top and Dress) and the rest is discarded. Datapoints with incomplete or missing features are removed as well as duplicate datapoints. In addition, the datasets were balanced. Features with a variance of less than 0.001 were eliminated. All the features were normalized (mean with 0 and variance with 1). To reduce the dimensionality of the features, all features were sorted according to their analysis of variance (ANOVA) F-value, and top 5 were selected. Different preprocessing routines are discussed in Appendix C.
@@ III-E Analysis
After all results from the experiments are agglomerated, the data is prepared for further analysis. The choice of the kernel function is treated as a hyperparameter. Our experiments have shown that the inner normalized and inner kernels lead to almost identical predictions and GDs as discussed in Section IV-B2, so we limit our main analysis to the inner and distance kernels. We therefore have only two kernel function which are encoded as the binary variable basis. We also exclude hyperparameter $C$ (as it has no influence on the GD) from the main analysis and discuss it separately in Section IV-B6. For datasets 1 and 8, we removed the highest 3% of each GD due to the presence of significant outliers. These outliers may occur when the classical kernel matrix has nearly singular values (Eq. 9) and the GD is therefore exceedingly large.
@@ IV Results
Fig. 1 visualizes the range of accuracies across different search grid settings as well as best achieved performances of classical and quantum learners (for more details on best learners see Appendix A). Since the quantum models achieve comparable results to the classical counterparts, we can assume that the explored hyperparameter ranges allowed for a high enough expressivity. Nevertheless, the best classical model can barely be surpassed. The selection of datasets and their characteristics range from those which are easy to learn, to those which are difficult to learn. There is also a varying spread of the accuracies on one specific dataset across the different parameter settings.
@@ IV-A Hyperparameter importance
We analyze whether accuracy and the GD metrics assign the same importance to the hyperparameters. For that we extract Gini importance from GBDT model fitted to the data extracted from the search grid experiment as described in Section III-E. Fig. 2 displays the average importance found for the different hyperparameters across all the datasets for performance as well as for the GD to the classical RBF kernel. See Fig. 14 for a breakdown of hyperparameter importance by all metrics.
@@
Judging from the results in the Fig. 2, the hyperparameter $t$ is assigned the most importance by both the accuracy and the GD metrics. The choice of the basis and $\gamma$ have a strong effect on these metrics as well, as both assigned these two parameters a similar importance that ranks them together at the second place. When examining the standard deviations, it is apparent that the importance of each hyperparameter for the GD is more consistent across the datasets than the importance for the accuracy.
@@ IV-B1 Evolution time ($t$)
Fig. 3 visualizes the marginals of the $t$ parameter for the accuracy and the GD. The standard deviations are illustrated as shadows and indicate the importance of $t$ on each of those metrics. If a standard deviation is small when averaging over all other parameters and datasets, this indicates that the metric does not depend on these other parameters as strongly as on $t$. This graphic highlights that the $t$ values for which a good performance was achieved corresponds to a rather small GD and as the predictions get worse the potential for quantum advantage increases. This behavior is consistent and sometimes even more significant when looking at individual datasets (Fig. 18 and Fig. 19) or 2-marginals like Fig. 3(a). Furthermore, of all hyperparameters, $t$ causes the most differences between the choices of basis. The distance kernel seems to be more resilient for smaller values of $t$, demonstrated in Fig. 17. As $t$ increases, the GD begins to rise earlier when the distance kernel is used.
@@
In order to have a fair comparison between the distance kernel and the inner kernel the $\gamma$ parameter is not averaged over, rather the maximal value over this parameter is chosen in all of Section IV, where the choice between the two is considered a parameter too (Section III-E). This is further motivated by the fact that $\gamma$ along with $C$ are the computationally cheapest hyperparameters to optimize, since the kernel can be reevaluated without having to recompute the distances between DMs. Fig. 5 shows the accuracies achieved for the specific choices of kernel function basis over all other parameter settings (except for $\gamma$) and datasets. If $\gamma$ is optimized, then the distance basis outperforms the other two. In Fig. 16, we observe a different situation when $\gamma$ is averaged like all the other parameters, but this is mostly due to the broad range (and thus many suboptimal values) we explored. When comparing the two normalization methods (inner and inner normalized) we notice almost no difference in performance or GD. This also applies to the best model for each dataset, as shown in Appendix A.

### v3

@@ Abstract
Quantum kernel methods are a promising method in quantum machine learning thanks to the guarantees connected to them. Their accessibility for analytic considerations also opens up the possibility of prescreening datasets based on their potential for a quantum advantage. To do so, earlier works developed the geometric difference, which can be understood as a closeness measure between two kernel-based machine learning approaches, most importantly between a quantum kernel and a classical kernel. This metric links the quantum and classical model complexities, and it was developed to bound generalization error. Therefore, it raises the question of how this metric behaves in an empirical setting. In this work, we investigate the effects of hyperparameter choice on the model performance and the generalization gap between classical and quantum kernels. The importance of hyperparameters is well known also for classical machine learning. Of special interest are hyperparameters associated with the quantum Hamiltonian evolution feature map, as well as the number of qubits to trace out before computing a projected quantum kernel. We conduct a thorough investigation of the hyperparameters across 11 datasets and we identify certain aspects that can be exploited. Analyzing the effects of certain hyperparameter settings on the empirical performance, as measured by cross validation accuracy, and generalization ability, as measured by geometric difference described above, brings us one step closer to understanding the potential of quantum kernel methods on classical datasets.
Index Terms:
Quantum machine learning, Quantum kernel methods, Hyperparameters, Geometric difference
@@ I Introduction
Assessing the performance of quantum models on current quantum computers in tackling machine learning tasks on classical datasets is an enduring ambition of the research community. Many strides have been made addressing this subject from theory-driven approaches that showed encouraging results for specific models, e.g., investigating the difference in capacity between classical and quantum models [2] and showcasing a model that allows to sample from a circuit that is hard to simulate classically [5], to more empirical-driven projects that showed promising results for specific models on specific use-cases, e.g., medical classification with hybrid quantum-classical convolutional neural network [15] and generation of high-resolution MNIST datapoints with hybrid generative neural networks [21]. However, making an overall claim of quantum advantage for a broader class of models for a broader set of use-cases remains challenging, especially due to the notorious insufficiency of theoretical understanding of certain sub-fields of machine learning, such as deep learning (DL) [12]. Nevertheless, in the recent years, a body of literature emerged linking DL models to kernel methods (KMs) [6, 9, 11] followed by literature making similar conclusion for the quantum domain [20, 8, 22]. This is great news due to favorable qualities of KMs, such as exhibiting convex loss landscapes, which allows to establish convergence, and hence trainability, guarantees [22]. It is a big advantage in comparison to DL that infamously suffer from non-convex loss landscapes and their quantum counterparts that suffers from barren plateaus [14]. Focusing on KMs therefore allows for a more theoretically sound argument, while at the same time, the insights gained for KMs can be applied to a broader range of models due to the aforementioned interrelationships.
@@
In this work, we strive to comprehend the capacity of quantum kernels on classical datasets. It is our ambition to understand the possible gap between classical and quantum models and whether GD can be a helpful ally for this purpose. However, given a rather theoretical nature of this metric, we need to first deepen our grasp of the empirical relevance of the GD tool. To this extent, we study how the choice of hyperparameters influences the GD and the empirical performance of the quantum KMs. We performed an experiment spanning multiple different classical datasets, a large hyperparameter search grid, and different kernel setups. Additionally, we artificially saturate the difference between classical and quantum learners for some datasets by creating artificial quantum label function following ideas in [7] to expand our understanding of GD metric. We empirically established which hyperparameters QKMs are most sensitive to and hence are the most crucial to optimize. Our results indicate that the GD depends on the individual hyperparameters as much as the accuracy, but the optimal values of them differ greatly between the metrics. These insights open new research revenues towards understanding which use-cases might be most well suited for QKMs.
@@ II-B Hamiltonian evolution feature map
For a quantum algorithm to operate on a classical data, the data has to be embedded into a quantum state $\phi:\mathbf{x}_{i}\to\ket{\mathbf{x}_{i}}$, which exists in a complex Hilbert space $\mathcal{H}$ that is $\mathbb{C}^{2^{n}}$ with $n$ being the number of qubits. The significance of the choice of embedding operation for the performance of quantum models has been highlighted in the literature [2, 23]. In this work, we concentrate on the Hamiltonian evolution feature map as it has been proven to be powerful in QKM setting [4, 7, 24]. This feature map has been developed for many-body problems, and it requires $n+1$ qubits to embed $n$ features and is defined as follows
@@
where $H_{j}^{XYZ}=(X_{j}X_{j+1}+Y_{j}Y_{j+1}+Z_{j}Z_{j+1})$ with $X_{j},Y_{j},Z_{j}$ being Pauli operators acting on $j$-th qubit, |$\ket{\psi_{j}}$ is a Haar-random state, and $t$ and $T$ are hyperparameters that symbolizes total evolution time and number of Trotter steps, respectively. The significance of hyperparameter $t$ for QKMs has been highlighted by Shaydulin et al. [24]. Our results support and extend their work.
@@ II-C Quantum kernels
Quantum kernels insert data into exponentially high dimensional Hilbert feature space, which is intractable for classical computers. The high dimensionality of this space has to be treated with caution as it requires $N\geq\Omega(2^{n})$ of training data to approximate a certain function well, where $n$ is the amount of logical qubits [7, 10, 19, 26]. This is due to a problem known as exponential concentration [26]. The same logic holds for a quantum fidelity kernel of mixed quantum states $k(x,x^{\prime})=Tr(\rho(x)\rho(x^{\prime}))$ [19]. Reducing the dimensionality of the space by projecting it onto a smaller subspace, thereby creating so-called projected quantum kernels, increases the generalization capacity of these methods. To achieve that, we can utilize reduced density matrices (RDMs) $\rho^{(q)}(x)=Tr_{\neq q}(\Phi(\rho(x)))$, where $Tr_{\neq q}$ stands for a trace of a subsystem that is created by excluding $q$ qubits and $\Phi$ is a trace-preserving map. Depending on the choice of the map $\Phi$, different proposals for projected kernels can be found in the literature [7, 10]. As generalized by Nakaji et al. [19], these projected (RDM-based) kernels can be categorized into inner product based defined as
@@ II-D Geometric Difference (GD)
Huang et al. [7] challenged a common argument for claiming quantum advantage that is often used in QML literature. In their work, they showed that a classical intractability of the function $f(x)$, that is represented by a quantum model, is not sufficient to claim a quantum advantage. Under certain circumstances, a classical learner $h(x)$ can learn to approximate $f(x)$ given a sufficient number $N$ of independent samples from a data distribution $\mathcal{D}$. They showed that an expected error between $f(x)$ and $h(x)$ can be upper bounded as follows:
@@
where $s_{K}(N)$ is the model complexity of the trained function $h(x)$ and $c>0$ is a constant. For that reason, with $N\propto s_{K}(N)/\epsilon^{2}$ datapoints a classical model can learn to predict the quantum function $f(x)$ up to an error $\epsilon$. Thus, if the best classical learner $h(x)$ has a low model complexity $s_{K}(N)$, it will not require too much data $N$ to learn the outputs of the quantum model $f(x)$. From here, it follows, that for a quantum advantage to exist it is required that there is the largest possible separation between complexities of quantum and classical models. Huang et al. [7] considered kernel models in their work and defined an asymmetric GD-based test between kernel matrices that captures this separation on a specific dataset. This test is defined as follows:
@@
where $K_{C}$ and $K_{Q}$ are classical and quantum kernel matrices with $Tr(K_{Q})=Tr(K_{C})=N$, respectively. The distance based kernel meets this requirement, while the inner product based kernel requires rescaling according to $\tilde{K}_{Q}=N\cdot K_{Q}/Tr(K_{Q})$ in the case of projected kernels, due to the fact that the diagonal of this kernel’s matrix contains purity related values (see Appendix A). Importantly, this test is reliant on the GD between embedded datapoints, without considering the labeling function. The GD now links the classical model complexity $s_{K_{C}}(N)$ and the quantum model complexity $s_{K_{Q}}(N)$ according to
@@ III Experimental setup
In this work we investigate how different parameter settings affect the empirical performance of quantum kernels, as well as their generalization ability, as determined by the GD. We have designed an experiment to evaluate a variety of different hyperparameter settings. In this section, we describe the parameters of the search grid used in our experiment. We varied the types of models and their hyperparameters and tested their performance on various datasets.
@@ III-A Models
In this work, we combine classical SVMs (see Section II) from sklearn library with quantum kernels implemented with PennyLane framework on simulators. To obtain the kernel, we first embed classical datapoints with the Hamiltonian evolution feature map as in Eq. 2 and the full DM of the resulting quantum state is saved. These DMs are used to compute RDM kernels described in Eq. 4 (with additional rescaling to $Tr(K_{Q})=N$), Eq. 5 and Eq. 9. From here forth, we will refer to these kernels as inner and distance kernels respectively. In all cases, we set $\alpha_{k}=1/N_{K}$. In Appendix F we discuss the case $\alpha_{k}=1$.
@@ III-C Datasets
To train and test our models, we selected the 11 different datasets listed in Table II that have been previously used in the QML literature [7, 16, 17, 24]. Due to simulation time restrictions, we compress each dataset’s feature space to 5 dimensions, which translates to 6 qubits according to Eq. 2, and 200 datapoints (or less features/datapoints if fewer were available). In this study, we restrict to binary classification tasks, so for multi-class dataset 8 (Fashion-MNIST) only two classes are selected (T-shirt/Top and Dress) and the rest are discarded. Datapoints with incomplete or missing features are removed as well as duplicate datapoints. In addition, the datasets were balanced. Features with a variance of less than 0.001 were eliminated. All the features were normalized (mean with 0 and variance with 1). To reduce the dimensionality of the features, all features were sorted according to their analysis of variance (ANOVA) F-value, and the top 5 were selected. Different preprocessing routines are discussed in Appendix D.
@@ III-E Analysis
After all results from the experiments are agglomerated, the data is prepared for further analysis. The choice of the kernel function is treated as a hyperparameter. The two kernel functions are encoded as the binary variable basis. We also exclude hyperparameter $C$ (as it has no influence on the GD) from the main analysis and discuss it separately in Section IV-B6. For datasets 1 and 8, we removed the highest 3% of each GD due to the presence of significant outliers. These outliers may occur when the classical kernel matrix has nearly singular values (Eq. 7) and the GD is therefore exceedingly large.
@@ IV Results
Fig. 1 visualizes the range of accuracies across different search grid settings as well as best achieved performances of classical and quantum learners (for more details on best learners see Appendix B). Since the quantum models achieve comparable results to the classical counterparts, we can assume that the explored hyperparameter ranges allowed for a high enough expressivity. Nevertheless, the best classical model can barely be surpassed. The selection of datasets and their characteristics range from those which are easy to learn, to those which are difficult to learn. There is also a varying spread of the accuracies on one specific dataset across the different parameter settings.
@@ IV-A Hyperparameter importance
We analyze whether accuracy and the GD metrics assign the same importance to the hyperparameters. For that, we extract Gini importance from the GBDT model fitted to the data extracted from the search grid experiment as described in Section III-E. Fig. 2 displays the average importance found for the different hyperparameters across all the datasets for performance as well as for the GD to the classical RBF kernel. See Fig. 15 for a breakdown of hyperparameter importance by all metrics.
@@
Judging from the results in Fig. 2, the hyperparameter $t$ is assigned the most importance by both the accuracy and the GD metrics. The choice of the basis and $\gamma$ has a strong effect on these metrics as well, as both assigned these two parameters a similar importance that ranks them together at the second place. When examining the standard deviations, it is apparent that the importance of each hyperparameter for the GD is more consistent across the datasets than the importance for the accuracy.
@@ IV-B1 Evolution time ($t$)
Fig. 3 visualizes the marginals of the $t$ parameter for the accuracy and the GD. The standard deviations are illustrated as shadows and indicate the importance of $t$ on each of those metrics. If a standard deviation is small when averaging over all other parameters and datasets, this indicates that the metric does not depend on these other parameters as strongly as on $t$. This graphic highlights that the $t$ values for which a good performance was achieved correspond to a rather small GD, and as the predictions get worse, the potential for quantum advantage increases. This behavior is consistent and sometimes even more significant when looking at individual datasets (Fig. 19 and Fig. 20) or 2-marginals like Fig. 4(a). Furthermore, of all hyperparameters, $t$ causes the most differences between the choices of basis. The distance kernel seems to be more resilient for smaller values of $t$, demonstrated in Fig. 18. As $t$ increases, the GD begins to rise earlier when the distance kernel is used.
@@
In order to have a fair comparison between the distance kernel and the inner kernel the $\gamma$ parameter is not averaged over, rather the maximal value over this parameter is chosen in all of Section IV, where the choice between the two is considered a parameter too (Section III-E). This is further motivated by the fact that $\gamma$ along with $C$ are the computationally cheapest hyperparameters to optimize, since the kernel can be reevaluated without having to recompute the distances between DMs. Fig. 5 shows the accuracies achieved for the specific choices of kernel function basis over all other parameter settings (except for $\gamma$) and datasets. If $\gamma$ is optimized, then the distance basis outperforms the other two. In Fig. 17, we observe a different situation when $\gamma$ is averaged like all the other parameters, but this is mostly due to the broad range (and thus many suboptimal values) we explored.